OpenStudy (anonymous):

pls help :) Each employee in a company that manufactures computer parts receives training sessions on stress management. According to the analysis of the company researcher, an employee can resemble N(x) computer parts in a day, N(x)=150x^2+30x / 17x^2+45 after x days of stress management training. Evaluate N'(4) and N'(6), then interpret this values.

OpenStudy (anonymous):

so for N(4) you have to plug in four wherever you see the value of x in the equations

OpenStudy (anonymous):

The values probably means the speed at which the employee can resemble N(4) and N(6) in a day.

OpenStudy (anonymous):

\[\frac{(50(4)^2+30x) }{ 17(4)^2+45 }\]

OpenStudy (anonymous):

so that is what it will look like when you set it up when it is N(4)

OpenStudy (anonymous):

It's 150...

OpenStudy (anonymous):

hey, dont give answers, we are here to help them work it out on their own

OpenStudy (anonymous):

Differentiate using the quotient rule \[\frac{ d }{ dx }(\frac{ u }{ v }) = \frac{ vu' - uv' }{v^2 }\]

OpenStudy (anonymous):

I meant your fraction... it's 150(4)^2...

OpenStudy (anonymous):

oh, ok, my bad. sorry lol.

OpenStudy (anonymous):

@sweetie21 are you following what we are saying?

OpenStudy (anonymous):

yea :)

OpenStudy (anonymous):

Just differentiate N with respect to x and substitute x = 4 and x = 6 into N'(x)